# Emergence of a High-Dimensional Abstraction Phase in Language Transformers

### v1

@@ 1 Introduction
Compression is thought to underlie generalizable representation learning (Deletang et al., 2024; Yu et al., 2023). Indeed, language models compress their input data to a manifold of dimension orders-of-magnitude lower than their embedding dimension (Cai et al., 2021; Cheng et al., 2023; Valeriani et al., 2023). Still, the intrinsic dimension (ID) of input representations may fluctuate over the course of processing: we ask, what does the evolution of ID over layers reveal about representational generalizability, and, broadly, about linguistic processing?
@@
Taken together, our experiments suggest that all analyzed transformer architectures develop, in the intermediate layers, a high-dimensional representation encoding complex and abstract linguistic information. The results of this processing are stored in representations which are then used, possibly through a process of incremental refinement, to predict the next token.
@@ 2 Related work
The notion that nominally complex, high-dimensional objects can be described using few degrees of freedom is not terribly new: it, for instance, underlies a number of popular dimensionality reduction methods such as PCA (Jolliffe, 1986). But, while PCA is linear, the data manifold need not be: as such, general nonlinear methods have been proposed to estimate the topological dimension, or manifold dimension, of point clouds (see Campadelli et al. (2015) for a survey). As neural representations tend to constitute nonlinear manifolds across modalities (Pope et al., 2021; Cai et al., 2021), we use a state-of-the-art nonlinear ID estimation method, the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022), which we further describe in Section 3.4.
Deep learning problems tend to be high-dimensional. But, recent work reveals that these ostensibly high-dimensional problems are governed by low-dimensional structure. It has, for instance, been shown that common learning objectives and natural image data lie on low-dimensional manifolds (Li et al., 2018; Pope et al., 2021; Psenka et al., 2024); that learning occurs in low-dimensional parameter subspaces (Aghajanyan et al., 2021; Zhang et al., 2023); that modern neural networks learn highly compressed representations of images, protein structure, and language (Ansuini et al., 2019; Valeriani et al., 2023; Cai et al., 2021); and moreover, that lower-ID tasks and datasets are easier to learn (Pope et al., 2021; Cheng et al., 2023).
@@
Complementary to parameter ID, a number of works focus on the ID of representations in LM activation space. In particular, Cai et al. (2021) were the first to identify low-dimensional manifolds in the contextual embedding space of (masked) LMs. Balestriero et al. (2023) linked representational ID to the scope of attention, and showed how toxicity attacks can exploit this relationship. Tulchinskii et al. (2023) demonstrated that representational ID can be used to differentiate human- and AI-generated texts. Cheng et al. (2023) established a relation between representational ID and information-theoretic compression, also showing that dataset-specific ID correlates with ease of fine-tuning. Closer to our aims, Valeriani et al. (2023) studied the evolution of ID across layers for vision and protein transformers, with a preliminary analysis of a single language transformer tested on a single dataset. Like us, they found that ID develops in different phases, with a consistent early peak followed by a valley, and less consistent markers of a second peak. We greatly extend their analysis of linguistic transformers by investigating the functional role of the main ID peak in five distinct LMs.
@@ 3.1 Models
All language models considered have between 6.5 and 8B parameters, 32 hidden layers, and a hidden dimension of $4096$. Moreover, they all inherit the architectural design of the decoder-only transformer (Vaswani et al., 2017) with a few notable changes: 1) layer normalization operations are applied before self-attention and MLP sublayers; 2) Pythia adds in-parallel the self-attention and MLP sublayer outputs to the residual stream, while in other models, self-attention outputs are added to the residual stream before the MLP; 3) Llama/Mistral/OLMo replace the ReLU activation function with SwiGLU (Shazeer, 2020), and Pythia instead uses GeLU (Hendrycks and Gimpel, 2016); 4) all models but OPT replace absolute positional embeddings with rotary positional embeddings (Su et al., 2022); 5) to facilitate self-attention computation, Llama uses grouped query attention (Ainslie et al., 2023), Mistral applies a sliding window attention, and Pythia adopts Flash Attention (Dao et al., 2022). In all cases, we use as our per-layer representations the vectors stored in the HuggingFace transformers library hidden__\__states variable. [^1] Each hidden__\__state vector corresponds to the representation in the residual stream(Elhage et al., 2021) after one attention and one MLP update.
In autoregressive models, an input sequence’s last token representation is the only one to contain information about the whole sequence. Furthermore, it is the only one that is decoded at the last layer to predict the next token. For these reasons, we choose to represent input sequences with their last token representation at each layer.
@@ 3.2 Data
Since we focus on model behavior in-distribution, we compute observables using three corpora that proxy models’ pre-training data (all accessed through HuggingFace): Bookcorpus (Zhu et al., 2015); the Pile (Gao et al. (2020); precisely, the Pile-10k subsample available on HuggingFace) and WikiText-103 (Merity et al., 2017). From each corpus, we randomly extract 5 non-overlapping partitions of 10k 20-token sequences (sequence length is counted according to each corpus tokenization scheme). [^2] We replicated the Pile-based Pythia and OPT experiments with sequences extended to 128 tokens. We obtained very similar results, confirming that the sequences we are using cover the typical contextual spans encode in model representations. We do not constrain the sequences to have any particular structure: in particular, their final element is not required to coincide with the end of a sentence. We additionally generated 5 partitions of 10k 20-token sequences from shuffled versions of the same corpora. The latter respect the source corpus unigram frequency distribution, but syntactic structure and semantic coherence are destroyed.
@@ 3.3 Probing and downstream tasks
To relate representations’ geometry to their content, we use the probing datasets of (Conneau et al., 2018), meant to capture the encoding of surface-related, syntactic and semantic information in LM representations. For each layer, we train a lightweight MLP probe from the hidden representation to each linguistic task. Details on tasks and training procedure are given in Appendix H.
If there is a relation between ID and semantic content, then ID may also predict ease-of-transfer to downstream NLP tasks. To test this claim, we consider two such tasks, sentiment classification of film reviews (Maas et al., 2011) and toxicity classification (Adams et al., 2017), for which we train binary linear classifiers on each hidden layer representation (training details in Appendix I).
@@ 3.4 Intrinsic Dimension
Real-world datasets tend to show a high degree of (possibly) non-linear correlations and constraints between their features (Tenenbaum et al., 2000). This means that, despite a very large embedding dimensionality, data typically lie on a manifold characterized by a much lower dimensionality, referred to as its intrinsic dimension (ID). This quantity may be thought of as the number of independent features needed to locally describe the data with minimal information loss.
In almost every real-world system, the ID depends on the scale at which the data is analysed. In particular, at small scales, the true dimensionality of the manifold is typically hidden by that of data noise. At very large scales, the ID can also be erroneous, due to, for instance, the curvature of the manifold (Facco et al., 2017; Denti et al., 2022). For this reason, in order to obtain a reliable and meaningful ID estimation, a proper scale analysis is necessary. To this aim, we opted for the generalized ratios intrinsic dimension estimator (GRIDE) of Denti et al. (2022), which extends the commonly used333(Ansuini et al., 2019; Valeriani et al., 2023; Tulchinskii et al., 2023; Cheng et al., 2023) TwoNN estimator of Facco et al. (2017) to general scales.
@@
For each (model, corpus, layer) combination, we perform an explicit scale analysis. To do so, we first estimate the ID while varying $k$. Then, by visual inspection, we select a suitable $k$ that coincides with a plateau in ID estimate (Denti et al., 2022). An example of such a scale analysis for one (model, corpus) combination is reported in Appendix C.
In order to delimit high-ID peaks across layers, we conventionally locate the end of the peak at the closest inflection point after its maximum value. The beginning of the peak corresponds, then, to the closest layer before the maximum with value equal or greater than that at the peak end.
@@ 3.5 Quantifying the relative information content of different representations
Similar to Valeriani et al. (2023), we wish to relate dimensional expansion or compression of representations across layers to changes in their neighborhood structure. If neighborhood structure defines a semantics in representational space (Boleda, 2020), then layers whose activations have similar neighborhood structures perform similar functions.
In particular, the layers of an LM are iterative reconfigurations of representation space. We quantify the extent of reconfiguration (conversely, stability) using a statistical measure called Information Imbalance (Glielmo et al., 2022), hereon referred to as $\Delta$. Given two different spaces $A$ and $B$, this quantity, defined in equation (1), measures the extent to which the neighborhood ranks in space $A$ are informative about the ranks in space $B$ (since ID computation is based on Euclidean distance, ranks are also obtained with Euclidean distance):
@@
It is essential to notice that $\Delta$ is non-commutative with respect to its arguments. That is, it is asymmetric upon a swap of spaces: $\Delta(A\rightarrow B)\neq\Delta(B\rightarrow A)$. This feature allows us to capture directional information containment, as we do below when considering the extent to which, e.g., a layer $i$ contains the information present in another layer $j$, or to measure the reciprocal information between representations of two different networks.
@@
In comparing the layers’ representation spaces, we also considered Doimo et al. (2020)’s neighborhood overlap measure which is used in Valeriani et al. (2023), as well as Representational Similarity Analysis (Kriegeskorte et al., 2008). These symmetric similarity measures tell a similar story to $\Delta$, but the signal is much weaker; therefore, we do not report them.
@@ 4.1 Emergence of a central high-dimensionality phase
Figure 1: Average ID (over 5 random partitions) of three corpora: Bookcorpus, the Pile and Wikitext, where samples come from (left) the original corpus, (right) the shuffled corpus. All curves are shown with $\pm$ 2 standard deviations (shuffled ID SDs are very small). In the middle layers, shuffled corpus ID is lower than non-shuffled ID, suggesting that linguistic processing contributes to ID expansion.
All models clearly go through a phase of high intrinsic dimensionality that tends to take place relatively early (starting approximately at layer 6 or 7, and mostly being over by layer 20), except for OPT, where it approximately lasts from layer 17 to layer 23. For all models, we also observe a second, less prominent peak occurring towards the end: only for OLMo, this second peak, which we do not analyze further, is as high as the first one.
@@
Figure 1 (right) shows that, when the models are fed shuffled corpora, ID remains low all throughout the layers. This suggests that the presence of high-ID peaks depends on the network performing meaningful linguistic processing. We further confirm this notion in Appendix D, where we analyze the evolution of the ID profile over the course of training for Pythia (the sole model whose intermediate checkpoints are public). We find that, over the course of training, the IDs’ magnitude not only grows over time, but that they become more peaked (Figure D.1), indicating that the characteristic profile emerges from learned structure.
@@
Figure 2 reports $\Delta(l_{i}\to l_{first})$ and $\Delta(l_{i}\to l_{last})$ for Llama, OPT and Pythia (with the remaining models in Appendix E). We observe that the ID peak largely overlaps with a peak in $\Delta(l_{i}\to l_{first})$. Here, a large value of $\Delta$ implies that sequences which are nearest neighbours at the ID-peak are distant from each other in the input layer. That is, the ID peak layers no longer contain the information encoded in the initial representation of the sequence, suggesting that they instead capture higher-level information. Meanwhile, we note that the $\Delta(l_{i}\to l_{last})$ profiles are not clearly related to the first ID peak, and we leave their analysis to future work.
@@
The high-dimensionality peaks contain similar information across models, as shown in the representative comparisons of Figure 4 (see Appendix G for the other pairs), which display cross-model $\Delta$ averaged across corpora and partitions. The high-ID layers always overlap with low-$\Delta$ layers, indicating that between models, representations at the peak have close neighborhood structures, and thus capture similar semantics.
Conversely, ID-peak layers have high cross-model $\Delta$ with layers outside the peak, shown by a lack of points outside the shaded intersection in Figure 4. This indicates that, while ID-peak representations predict those of other models, they contain different information from other models’ representations outside of the peak.
We notice, moreover, a marked asymmetry in which the ID-peak layers of Pythia and OPT, the two models with the highest absolute IDs, directionally contain other models’ representations. On the other hand, when models have similar maximum IDs (such as in the Pythia vs. OPT comparison), $\Delta$ is more symmetric and very small, implying that their representation spaces are really equivalent.

### v2

@@ 1 Introduction
Compression is thought to underlie generalizable representation learning (Deletang et al., 2024; Yu et al., 2023). Indeed, language models compress their input data to a manifold of dimension orders-of-magnitude lower than their embedding dimension (Cai et al., 2021; Cheng et al., 2023; Valeriani et al., 2023). Still, the intrinsic dimension (ID) of input representations may fluctuate over the course of processing. We ask: what does the evolution of ID over layers reveal about representational generalizability, and, broadly, about linguistic processing?
@@
Taken together, our experiments suggest that all analyzed transformer architectures develop, in the intermediate layers, a high-dimensional representation encoding complex and abstract linguistic information. The results of this processing are stored in representations which are then used, possibly through a process of incremental refinement, to predict the next token. Besides providing new insights on the inner workings of Transformer-based LMs, these results have implications for tasks such as layer-based fine-tuning, model pruning and model stitching.
@@ 2 Related work
The notion that nominally complex, high-dimensional objects can be described using few degrees of freedom is not new: it underlies, for instance, a number of popular dimensionality reduction methods such as standard PCA (Jolliffe, 1986). But, while PCA is linear, the data manifold need not be: as such, general nonlinear methods have been proposed to estimate the topological dimension, or manifold dimension, of point clouds (see Campadelli et al. (2015) for a survey). As neural representations tend to constitute nonlinear manifolds across modalities (Ansuini et al., 2019; Cai et al., 2021), we use a state-of-the-art nonlinear ID estimation method, the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022), which we further describe in Section 3.4.
Deep learning problems tend to be high-dimensional. But, recent work reveals that these ostensibly high-dimensional problems are governed by low-dimensional structure. It has, for instance, been shown that common learning objectives and natural image data lie on low-dimensional manifolds (Li et al., 2018; Pope et al., 2021; Psenka et al., 2024); that learning occurs in low-dimensional parameter subspaces (Aghajanyan et al., 2021; Zhang et al., 2023); that modern neural networks learn highly compressed representations of images, protein structure, and language (Ansuini et al., 2019; Valeriani et al., 2023; Cai et al., 2021); and, moreover, that lower-ID tasks and datasets are easier to learn (Pope et al., 2021; Cheng et al., 2023).
@@
Complementary to parameter ID, a number of works focus on the ID of representations in LM activation space. In particular, Cai et al. (2021) were the first to identify low-dimensional manifolds in the contextual embedding space of (masked) LMs. Balestriero et al. (2023) linked representational ID to the scope of attention and showed how toxicity attacks can exploit this relationship. Tulchinskii et al. (2023) demonstrated that representational ID can be used to differentiate human- and AI-generated texts. Cheng et al. (2023) established a relation between representational ID and information-theoretic compression, also showing that dataset-specific ID correlates with ease of fine-tuning.
Closer to our aims, Valeriani et al. (2023) studied the evolution of ID across layers for vision and protein transformers, with a preliminary analysis of a single language transformer tested on a single dataset. Like us, they found that ID develops in different phases, with a consistent early peak followed by a valley and less consistent markers of a second peak. We greatly extend their analysis of linguistic transformers by investigating the functional role of the main ID peak in five distinct LMs and a number of new analyses and experiments. Importantly, our converging evidence suggests that, in language transformers, semantic processing of the input first takes place under the first ID peak, contra the preliminary evidence by Valeriani and colleagues that this crucial phase is the low-ID “valley” between the peaks.
@@ 3.1 Models
All language models considered have between 6.5 and 8B parameters, $32$ hidden layers, and a hidden dimension of $4096$. Moreover, they all inherit the architectural design of the decoder-only transformer (Vaswani et al., 2017) with a few notable changes: 1) layer normalization operations are applied before self-attention and MLP sublayers; 2) Pythia adds the self-attention and MLP sublayer outputs to the residual stream in parallel, while in other models self-attention outputs are added to the residual stream before the MLP; 3) Llama/Mistral/OLMo replace the ReLU activation function with SwiGLU (Shazeer, 2020), and Pythia uses GeLU instead (Hendrycks & Gimpel, 2016); 4) all models but OPT replace absolute positional embeddings with rotary positional embeddings (Su et al., 2022); 5) to facilitate self-attention computation, Llama uses grouped query attention (Ainslie et al., 2023), Mistral applies a sliding window attention, and Pythia adopts Flash Attention (Dao et al., 2022). In all cases, we use as our per-layer representations the vectors stored in the HuggingFace transformers library hidden__\__states variable. [^1] Each hidden__\__state vector corresponds to the representation in the residual stream(Elhage et al., 2021) after one attention and one MLP update.
In autoregressive models, due to the causal nature of attention, an input sequence’s last token representation is the only one to contain information about the whole sequence. Furthermore, it is the only one that is decoded at the last layer to predict the next token. For these reasons, we choose to represent input sequences with their last token representation at each layer.
@@ 3.2 Data
Since we focus on model behavior in-distribution, we compute observables using three corpora that proxy models’ pre-training data (all accessed through HuggingFace): Bookcorpus (Zhu et al., 2015); the Pile (Gao et al. (2020); precisely, the 10k-document subsample available on HuggingFace) and WikiText-103 (Merity et al., 2017). From each corpus, we randomly extract 5 non-overlapping partitions of 10k 20-token sequences (sequence length is counted according to the number of typographic tokens in the text of each corpus, as made available on HuggingFace). [^2] We replicated the Pile-based Pythia and OPT experiments with sequences extended to 128 tokens. We obtained very similar results, confirming that the sequences we are using cover the typical contextual spans encoded in model representations. We do not constrain the sequences to have any particular structure: in particular, their final element is not required to coincide with the end of a sentence. We additionally generated 5 partitions of 10k 20-token sequences from shuffled versions of the same corpora. The latter respect the source corpus unigram frequency distribution, but syntactic structure and semantic coherence are destroyed.
@@ 3.3 Probing and downstream tasks
To relate representations’ geometry to their content, we use the probing datasets of Conneau et al. (2018), meant to capture the encoding of surface-related, syntactic, and semantic information in LM representations. For each layer, we train a lightweight MLP probe from the hidden representation to each linguistic task. Details on tasks and training procedure are given in Appendix I.
If there is a relation between ID and abstract linguistic processing, then ID may also predict ease-of-transfer to downstream NLP tasks. To test this claim, we consider two such tasks in a zero-shot setting, sentiment classification of film reviews (Maas et al., 2011) and toxicity classification (Adams et al., 2017), for which we train binary linear classifiers on each hidden layer representation (training details in Appendix J). Note that we do not fine-tune the layers themselves, as this might change their IDs, which is the very property we want to relate to task performance.
@@ 3.4 Intrinsic Dimension
Real-world datasets tend to show a high degree of possibly non-linear correlations and constraints between their features (Tenenbaum et al., 2000). This means that, despite a very large embedding dimension, data typically lie on a (locally smooth) manifold characterized by a much lower dimensionality, referred to as its intrinsic dimension (ID). This quantity may be thought of as the number of independent features needed to locally describe the data with minimal information loss (Bishop, 1995). Equivalently, it has been defined as the dimensionality of the support of the probability distribution from which the data is generated (Fukunaga, 2013; Campadelli et al., 2015).
In almost every real-world system, the ID depends on the scale, i.e., the size of neighbourhood at which the data is analyzed. In particular, at small scales, the true dimensionality of the manifold is typically hidden by that of data noise. At very large scales, the ID estimate can also be erroneous, due, for instance, to the curvature of the manifold (Facco et al., 2017; Denti et al., 2022). For this reason, in order to obtain a reliable and meaningful ID estimation, a proper scale analysis is necessary. This is typically performed by varying the amount of the neighbours considered in estimating the ID and looking for an interval of scales in which the estimate is approximately stable (see Appendix C for further details). To this aim, we opted for the generalized ratios intrinsic dimension estimator (GRIDE) of Denti et al. (2022), which extends the commonly used [^3] E.g., Ansuini et al. (2019); Tulchinskii et al. (2023); Valeriani et al. (2023). TwoNN estimator of Facco et al. (2017) to general scales.
@@
For each (model, corpus, layer) combination, we perform an explicit scale analysis. To do so, we first estimate the ID while varying $k$. Then, by visual inspection, we select a suitable $k$ that coincides with a plateau in ID estimate (Denti et al., 2022). An example of such a scale analysis is reported in Appendix C, where we also demonstrate our estimates to be robust to changes in scale.
In order to delimit high-ID peaks across layers, we conventionally locate the end of the peak at the closest inflection point after its maximum value. The beginning of the peak corresponds, then, to the last layer before the maximum with value equal or greater than that at the end of the peak.
@@ 3.5 Quantifying the relative information content of different representations
We wish to relate dimensional expansion or compression of representations across layers to changes in their neighborhood structure. If neighborhood structure defines a semantics in representational space (Boleda, 2020), then layers whose activations have similar neighborhood structures perform similar functions.
In particular, the layers of an LM are iterative reconfigurations of representation space. We quantify the extent of reconfiguration (conversely, stability) using a statistical measure called Information Imbalance (Glielmo et al., 2022), hereon referred to as $\Delta$. Given two different spaces $A$ and $B$, this quantity, defined in equation 1, measures the extent to which the neighborhood ranks in space $A$ are informative about the ranks in space $B$ (since ID computation is based on Euclidean distance, ranks are also obtained with Euclidean distance):
@@
In comparing the layers’ representation spaces, we also considered three alternative measures: Doimo et al. (2020)’s neighborhood overlap measure, also used by Valeriani et al. (2023), Representational Similarity Analysis (Kriegeskorte et al., 2008), and linear CKA (Kornblith et al., 2019). Only the last measure resulted in interpretable results, broadly coherent with those obtained with $\Delta$ (Appendix H).
We choose to focus on $\Delta$ as the most principled measure for the investigation of LM layers, where we can make very little assumptions about the shape of the manifold. Moreover, crucially, $\Delta$, unlike the other measures we considered, is non-commutative with respect to its arguments. That is, it is asymmetric upon a swap of spaces: $\Delta(A\rightarrow B)\neq\Delta(B\rightarrow A)$. This feature allows us to capture directional information containment. For example, we will see below that $\Delta$ detects the asymmetric relation between Pythia/OPT and Lllama representations (Figure 4), and it allows us to study the degree to which information in a certain layer directionally predicts information in the first, next, or last layer (Figure 2 and Figure 3).
@@ 4.1 Emergence of a central high-dimensionality phase
Figure 1: Average ID over layers over 5 random partitions from each of three corpora: Bookcorpus, the Pile and Wikitext. (Left): different models’ layerwise ID plotted for the original corpora. (Center): different models’ layerwise ID plotted for the shuffled corpora. (Right): different Pythia training checkpoints’ layerwise ID on the original corpora, where darker curves are later checkpoints. In the middle layers, shuffled corpus ID (center) is lower than non-shuffled ID (left), suggesting that linguistic processing contributes to ID expansion. ID increases over the course of training for all layers to reach the final profile at step 143000 (right), suggesting that ID reflects learned linguistic features. All curves are shown with $\pm$ 2 standard deviations (shuffled SDs are very small).
All models clearly go through a phase of high intrinsic dimensionality that tends to take place relatively early (starting approximately at layer 6 or 7, and mostly being over by layer 20), except for OPT, where it approximately lasts from layer 17 to layer 23. For all models, we also observe a second, less prominent peak occurring towards the end: only for OLMo, this second peak, which we do not analyze further, is as high as the first one. Appendix D reports ID profiles for a smaller and a larger LM from the Pythia family, showing that the presence of the ID peak is not dependent on the specific size range we are focusing on.
@@
Figure 1 (center) shows that, when the models are fed shuffled corpora, ID remains low all throughout the layers. This suggests that the presence of high-ID peaks depends on the network performing meaningful linguistic processing. We further confirm this notion in Figure 1 (right), where we analyze the evolution of the ID profile over the course of training for Pythia (whose intermediate checkpoints are public). We find that, over the course of training, the IDs’ magnitude not only grows over time, but that they become more peaked, indicating that the characteristic profile emerges from learned structure.
@@
Figure 2 reports $\Delta(l_{i}\to l_{first})$ and $\Delta(l_{i}\to l_{last})$ for Llama, OPT and Pythia (with the remaining models in Appendix E). We observe that the ID peak largely overlaps with a peak in $\Delta(l_{i}\to l_{first})$. Here, a large value of $\Delta$ implies that sequences which are nearest neighbours at the ID-peak are distant from each other in the input layer. That is, the ID peak layers no longer contain the information encoded in the initial representation of the sequence, suggesting they instead capture higher-level information. Meanwhile, we note that the $\Delta(l_{i}\to l_{last})$ profiles are not clearly related to the first ID peak, and we leave their analysis to future work.
@@
The high-dimensionality peaks contain similar information across models, as shown in the representative comparisons of Figure 4, which display cross-model $\Delta$ averaged across corpora and partitions (see Appendix G for the other pairs, and Appendix H for broadly comparable patterns emerging from CKA). The high-ID layer intersections always correspond to regions with a concentration of low-$\Delta$ values, indicating that, between models, representations at the peak have close neighborhood structures, and thus capture similar semantics (note, still, that low-$\Delta$ values can also occur outside the peak intersections, suggesting high ID might be a sufficient, rather than necessary, condition for cross-model similarity).
We also notice a marked asymmetry in which the ID-peak layers of Pythia and OPT, the two models with the highest absolute IDs, directionally contain other models’ representations. On the other hand, when models have similar maximum IDs (such as in the Pythia vs. OPT comparison), $\Delta$ is more symmetric and very small, implying that their representation spaces are really equivalent.